A block of mass 5.0 kg is suspended from the end of a vertical spring which is stretched by 10 cm under the load of the block. The block is given a sharp impulse from below so that it acquires an upward speed of 2.0 m/s. How high will it rise? Take g = 10 m/s2.

Grade:10

1 Answers

Deepak Patra
9 years ago
Sol. m = 5kg, x = 10cm = 0.1m, v = 2m/sec, h =? G = 10m/sec^2 S0, k = mg/x = 50/0.1 = 500 N/m Total energy just after the blow E = ½ mv^2 + ½ kx^2 …(i) Total energy a a height h = ½ k (h - x)^2 + mgh …..(ii) ½ mv^2 + ½ kx^2 = ½ k (h - x)^2 + mgh On, solving we can get, H = 0.2 m = 20 cm